### v1

# Optimizing Product Deduplication in E-Commerce with Multimodal Embeddings

Abstract
In large scale e-commerce marketplaces, duplicate product listings frequently cause consumer confusion and operational inefficiencies, degrading trust on the platform and increasing costs. Traditional keyword-based search methodologies falter in accurately identifying duplicates due to their reliance on exact textual matches, neglecting semantic similarities inherent in product titles. To address these challenges, we introduce a scalable, multimodal product deduplication designed specifically for the e-commerce domain. Our approach employs a domain-specific text model grounded in BERT architecture in conjunction with MaskedAutoEncoders for image representations. Both of these architectures are augmented with dimensionality reduction techniques to produce compact 128-dimensional embeddings without significant information loss. Complementing this, we also developed a novel decider model that leverages both text and image vectors. By integrating these feature extraction mechanisms with Milvus, an optimized vector database, our system can facilitate efficient and high-precision similarity searches across extensive product catalogs exceeding 200 million items with just 100GB of system RAM consumption. Empirical evaluations demonstrate that our matching system achieves a macro-average F1 score of 0.90, outperforming third-party solutions which attain an F1 score of 0.83. Our findings show the potential of combining domain-specific adaptations with state-of-the-art machine learning techniques to mitigate duplicate listings in large-scale e-commerce environments.
§§footnotetext: Equal contribution
1 Introduction
In today’s vast e-commerce marketplaces, particularly within the Turkish e-commerce landscape, customers frequently encounter duplicate product listings that create confusion and frustration during shopping. These duplicates make it difficult to compare options, slow down decision-making, and diminish trust in the platform. From an operational standpoint, maintaining a catalog with duplicate products creates significant overhead—increased storage and computational costs for search systems. Moreover, duplicate listings can disrupt Buy Box management, a crucial feature that determines which seller’s product is highlighted as the default option for purchase. When duplicates exist, multiple sellers offering the same product might unintentionally compete against themselves, splitting sales metrics and reducing their likelihood of winning the Buy Box. This can lead to unnecessary price wars and complicate marketplace analytics. Effective product deduplication ensures a clean catalog, allowing the Buy Box to function optimally by consolidating seller competition and attributing sales accurately. As a result, it enhances both operational efficiency and customer satisfaction.
The challenge of product deduplication grows with the scale and diversity of product listings. In the Turkish e-commerce sector, this problem is particularly acute due to the rapid growth of the market and the high volume of unique sellers. Traditional keyword-based search methods fail to effectively identify duplicates because they rely on exact text matches rather than semantic similarities. For example, ”wireless earbuds” and ”Bluetooth earphones” might describe identical products but would be treated as different items by keyword-based systems. This limitation has led to the adoption of vector search-based solutions, which represent products as high-dimensional vectors for semantic similarity comparisons. While vector search methods excel at capturing nuanced relationships in product data, deploying such systems at scale in e-commerce introduces its own set of challenges, including inference latency, memory overhead, and maintaining efficiency in a growing catalog dataset with over 220 million products.
Existing solutions often leverage pretrained models due to their strong capabilities in generating meaningful representations. However, these models produce high-dimensional vectors. For example, our initial implementation using the Google Universal Sentence Encoder Multilingual [1] model generated 512-dimensional embeddings. While effective, it faced scalability issues when we hit indexing 8 million items, causing inference times to rise and operational costs to increase. Such limitations highlight the need for scalable, efficient models tailored to the specific requirements of e-commerce.
To address these bottlenecks in the vector search capabilities, we developed a domain-specific text model, fine-tuned for the linguistic and product-specific nuances of Turkish e-commerce, which generates compact product representation. The model built on BERT [2], incorporating dimensionality reduction and enhanced feature extraction through intermediate transformer layers and convolutional layers. This approach enables us to compress representations to 128 dimensions without significant information loss.
Additionally, we designed a decider model to classify product pairs as duplicates, leveraging text and image vectors to produce confidence scores. This classification-based architecture avoids the inefficiencies of pairwise comparisons inherent in siamese or contrastive learning, ensuring real-time inference and category-agnostic scalability. For the image representations that the decider model is used, we transitioned from traditional AutoEncoders to Masked AutoEncoders (MAE) with a structured patch selection logic, ensuring the preservation of critical visual details unique to e-commerce product images.
By integrating these advancements, our solution achieves improved scalability, efficiency, and accuracy in product deduplication. The proposed system addresses the unique challenges of the e-commerce domain, offering significant operational benefits and enhancing the customer shopping experience.
2 Related Works
This section presents a review of related works on feature extraction from text and image data, with a specific focus on applications in the e-commerce domain. We also inspect established approaches to product unification.
2.1 Text Based Feature Extraction
Transformer-based architectures have revolutionized semantic representation techniques through sophisticated contextual embedding models. For the Turkish language, there are notable publicly available pre-trained models which have been specifically trained on Turkish text with transformer-based architectures such as BERT, DistilBERT, and ConvBERT [3, 2, 4, 5], each offering unique advantages in terms of performance and efficiency.
BERT (Bidirectional Encoder Representations from Transformers) captures deep bidirectional context by considering both left and right sentence contexts simultaneously [2]. DistilBERT compresses BERT through knowledge distillation, reducing complexity while retaining comparable performance [4]. ConvBERT enhances local feature extraction and efficiency by integrating span-based dynamic convolutions [5].
Domain-specific language model adaptation has emerged as a critical research area in NLP. Gururangan et al. [6] demonstrated the effectiveness of continuing pretraining on domain-specific data before fine-tuning on labeled tasks. Ethayarajh [7] revealed that higher transformer layers capture increasingly context-dependent semantic features.
2.2 Image Based Feature Extraction
Generating representative feature vectors is an important step in image understanding. Convolutional Neural Networks (CNNs) are a cornerstone field in computer vision and paved the way of extracting image features without relying on hand crafted features [8]. Learned image features can be used in various applications in e-commerce such as image search [9, 10], product recommendation [11].
In order to extract image vectors by capturing maximum possible contextual image representation without an extensive labeling process, unsupervised learning based feature extraction systems are developed [12]. Extracted features on these models result in a large vector which is not suitable for deployment in the real world.
2.3 Foundational Multimodal Models
The rapid advancement of vision-language pre-training (VLP) has led to the development of powerful foundational models that learn joint representations of text and images. Contrastive Language-Image Pre-training (CLIP) is a prominent example, employing a dual-encoder architecture to align a vision transformer and a text transformer in a shared latent space via a contrastive loss. This approach, trained on a massive 400 million image-text dataset, enables zero-shot classification and cross-modal retrieval by learning a broad, holistic understanding of visual and textual concepts [13]. While highly general, CLIP’s primary focus on image-level understanding can be a limitation for fine-grained tasks, a challenge that has been addressed by subsequent works such as Alpha-CLIP [14], which introduces an auxiliary channel for region-specific attention, and Long-CLIP [15], which extends the text encoder’s token limit.
Building upon this, BLIP-2 introduced an efficient two-stage bootstrapping strategy that leverages frozen, pre-trained image encoders (e.g., from CLIP) and frozen Large Language Models (LLMs) [16]. This method uses a lightweight Querying Transformer (Q-Former) to bridge the modality gap, achieving state-of-the-art performance on tasks like visual question answering with significantly fewer trainable parameters than previous methods. While these models excel at general-purpose tasks, their architectural and training philosophies differ from our domain-specific approach, which is meticulously designed for the operational realities of e-commerce.
2.4 Product Deduplication Approaches
Matching similar products is an important part of a catalog management system in e-commerce. Duplicate products bloat the product catalog and may cause issues on both operational side and user experience. On the operations side, it increases costs due to increased or redundant manual labor for each product listing. On the other hand, users face redundant search results and overall poor user experience. In order to alleviate these issues,various systems [17, 18, 19, 20, 21] are proposed.
3 Theoretical Analysis
In the e-commerce domain, product deduplication is a critical component of catalog management. Traditional keyword-based approaches often fall short due to inconsistencies in product titles and attribute usage. To address this, we implemented a vector search-based deduplication pipeline that generates robust representations of products using both textual and visual data.
The vector search retrieves the top-N candidate products using text-based vectors. However, identifying true duplicates from among these candidates proved non-trivial. Early attempts to rely on vector distances and rule-based filtering were insufficient due to the diverse ways sellers create product listings. This necessitated a dedicated classification model—referred to as the decider model—that could learn to identify duplicates using the semantic and visual context of the products.
While product images are essential for accurate comparison, generating visual vectors for all catalog items is computationally expensive. To optimize infrastructure use, image vectors were only employed in the decider model, whereas the initial vector search operated solely on text-based representations.
4 Methodology
Our multimodal product matching system was developed through several iterations, focusing on robust feature extraction, efficient dimensionality reduction, and an accurate decision model. This section details the individual steps that resulted in our final architecture.
4.1 Feature Extraction and Dimensionality Reduction
A central challenge in deploying deep models at scale is managing the high dimensionality of representation vectors. Pretrained models like BERTurk [3] and EfficientNetV2 [22] yield 768- and 1792-dimensional vectors, respectively. These vectors are expensive to store, index, and search in large catalogs.
To determine the extent of feasible dimensionality reduction, we conducted a feasibility study. A binary decision model was trained using original and compressed versions of BERTurk and EfficientNetV2 representations. The original configuration achieved an F1 score of 0.86. Reducing both embeddings to 512 dimensions using PCA slightly changed the performance to 0.856. Further compression to 256 and 128 dimensions yielded F1 scores of 0.85 and 0.84, respectively. These results demonstrated that deduplication could be performed effectively using 128-dimensional vectors, which were adopted for all subsequent experiments to ensure scalability and efficiency.
Reducing dimensionality without compromising representation quality required more than simple projection. For textual data, we experimented with using token embeddings from intermediate BERT layers, which strike a balance between raw and highly contextualized representations. To further enhance efficiency, we added a CNN-based aggregation layer to automatically extract salient features from intermediate layers, improving performance while controlling vector size.
A similar approach was taken for images. Our initial solution involved using representations from an EfficientNet-based autoencoder. Encouraged by results from textual modeling, we later trained Masked AutoEncoders (MAEs), which preserve object-level contextual cues rather than superficial visual features. The initial approach of the model adapted only output dimensionality; the latest version employs guided patch selection to retain important visual details in the compressed representations.
4.1.1 Text Feature Models
We evaluated several open-source models to generate initial text representations. Among Turkish pretrained language models, as shown in Table I, BERTurk achieved the best performance on our masked language modeling (MLM) task, outperforming ConvBERTurk and DistilBERTurk [3].
TABLE I: Text Model Variants MLM Validation Scores
Model Name Validation Loss
bertTurk 1.51
convbertTurk 1.89
distillbertTurk 1.81
The initial text representation model employed BERT as a backbone, followed by a dimensionality reduction module (linear + regularization + activation) for compression and a classification module (linear + activation) for training. After training, the compressed “CLS” token served as the input sentence representation.
Building on this, a stricter feature extraction approach was adopted while maintaining the fixed 128-dimensional constraint. To achieve this, a compression algorithm was designed to refine the “CLS” token representation. Since using the “CLS” token in an MLM framework required a decoder, a simple restacking and shaping scheme was introduced.
Figure 1: Updated version of text model
To enhance information density and capture lower-level text features, intermediate BERT layer outputs are combined and passed to the compression step (see Figure 1). Non-consecutive intermediate layers were selected, stacked, and processed through convolutional layers before being incorporated similarly to the original BERT output. This modification enabled the model to leverage early Transformer representations, improving its ability to extract richer e-commerce text features.
4.1.2 Visual Feature Models
For the image representations, DeiT [23], Swin Transformer [24], and LeViT [25] were benchmarked on the masked autoencoding (MAE) task. DeiT achieved the highest Multi-scale Structural Similarity Index Measure (MS-SSIM) [26], followed by Swin and LeViT. Results are shared in Table II.
TABLE II: Image Model Variants MAE Validation Scores
Model MS-SSIM
LeViT 0.27
Swin Transformer 0.43
DeiT 0.45
The initial approach for dense image representation extraction began with an AutoEncoder using an EfficientNetV2 encoder and a convolutional decoder. To maximize information density, no skip connections between encoder and decoder were made; the decoder relied solely on the compressed 128-length vector. However, analysis revealed a lack of contextual understanding: while similar objects in size and color had close representations, variants of the same object were mapped to distant vectors.
Exploring contextual image understanding, DeiT-based regular AutoEncoders were tested, but concerns over similar pitfalls led to adopting a Masked AutoEncoder (MAE) approach. Unlike traditional AutoEncoders, MAEs reconstruct masked image regions from unmasked patches, enforcing contextual learning (see Figure 2-a).
In detail, standard MAEs divide images into “N” patches, mask a portion, and reconstruct the missing areas. Given our dimensional constraints, each unmasked patch was compressed to 32-length vectors, forming a total representation of 128. However, this traditional masking approach risked losing fine-grained details crucial in e-commerce images, such as product prints or brand markings. Additionally, as most image information is centrally located, randomly masking could result in samples missing key details.
To address this, we introduced a structured patch selection method (see Figure 2-b):
  • Divide the image into “k” equal patches.
  • Always select the center patch (containing the most information).
  • Select two random patches from the remaining eight.
  • Resize the full image to match a single patch size and use as the last patch.
  • Stack patches as standard MAE inputs.
Figure 2: (a) Second iteration of the image model P# denotes the selected patches, M# denotes the masked patches (b) Final iteration of the image model P# denotes the selected patches, C is the center patch, Xr is resized image
This approach preserved some randomness while ensuring key details were captured. Unlike traditional MAEs predicting only masked regions, our model reconstructed the full image, improving its ability to differentiate subtle variations.
4.1.3 Domain Adaptation
Since pretrained models were trained on general-purpose datasets, domain adaptation was crucial. We investigated both supervised and unsupervised methods. Supervised training on leaf-category prediction was found to encourage excessive clustering within product categories, potentially masking distinctions between similar items (e.g., t-shirt and skirt in a single image).
Unsupervised approaches proved to be more effective. In both cases, the goal was to reconstruct the original inputs from compressed vectors, thereby forcing the models to preserve semantically important features. This reconstruction-based training helped improve the quality of the representations without relying on noisy or synthetic labels.
4.2 Product Matching Pipeline
Storing, continuously adding new, and updating product vectors necessitate the use of a dedicated vector database (VecDB) that ensures high availability. Conducting searches using a brute-force approach is resource-intensive; thus, most contemporary VecDB solutions incorporate faster approximate nearest neighbor (ANN) search capabilities through various indexing algorithms [27].
For the VecDB setup, Milvus [28] was chosen for its active open-source community support, ease of setup, distributed capabilities, and its ability to execute various indexing and searching algorithms. The selection of indexing algorithms is crucial for optimizing the efficiency and speed of similarity searches. The two primary candidates considered were the clustering-based Inverted File FLAT (IVF_FLAT) [29] and the graph-based Hierarchical Navigable Small World (HNSW) [30]. Quantized indexing algorithms were deliberately excluded to prevent unnecessary loss in recall rate. Initial tests with varying vector counts and dimensions across both algorithms indicated that, given our memory constraints, IVF_FLAT emerged as the more suitable approach, provided that the appropriate cluster count and query units parameters are selected.
Development requirements for the decider model arose from the need to quickly determine if a pair of product listings belong to the same product. Initial trials using only vector distances showed promise; however, varying threshold values across different categories complicated management. To address this, a separate model was developed, leveraging dense vector representations to decide if given vectors correspond to the same product.
The decider model processes four vectors—two for image and two for text—associated with each product listing and generates a confidence score indicating the likelihood of them being the same product. A key architectural decision was to implement the model as a traditional classifier rather than using siamese or contrastive learning approaches. This choice was motivated by the need for efficient real-time inference, as standard classifiers can rapidly produce confidence scores from stored vectors. In contrast, siamese networks require computationally intensive pairwise comparisons, which are less scalable for large datasets.
Additionally, data availability and quality influenced this decision. Siamese and contrastive learning models typically rely on both paired and unpaired data to learn meaningful embeddings. However, only paired data was readily available, and generating synthetic unpaired data risked oversimplifying relationships, potentially hindering the model’s ability to learn robust distinctions. Moreover, a classifier inherently handles multiple product categories within a unified framework, eliminating the need for category-specific thresholds and simplifying system scalability and management.
By choosing a conventional classification approach, the decider model achieves efficient inference, better aligns with available data, and reduces operational complexity. This ensures high performance and robustness across diverse product categories, making the model well-suited for real-world applications in product listing management. The final system design can be seen in Figure 3.
Figure 3: Final system design for product deduplication.
5 Experiments & Results
Training of the architecture utilized a curated e-commerce dataset derived from internal product listings, consisting of 13 million products across 4,500 different categories, each represented by its title and the first product image. With a brief study on the data, for the text representations, the number of max tokens was determined to be 32, which covered %95 of the collected data. In addition, it was apparent that for an e-commerce listing title, most important words tend to exist at the beginning of sentences. Which made taking the first 32 tokens of a product title, a feasible approach.
From the 32 tokens fed as input, BERT created $32\times 128$ length representations with the help of a linear layer for dimensionality reduction. Afterwards for each of the masked input tokens, a classifier layer learned to predict the actual token. After the conclusion of the training, CLS token for each sentence was used as the final text representation. The base model trained on the masked language modeling (MLM) task achieved a validation loss of 1.51.
After further examination of the results and requirements regarding the product titles, we recognized that while BERT effectively captures contextual information within sentences, supplementing its representation vector with lower-level information could lead to additional improvements. To achieve this, intermediate transformer layer outputs from BERT—specifically from layers 1, 3, 5, 7, 9, and 11—were extracted alongside the outputs from the last transformer layer. These layers were then stacked and passed through a convolutional chain followed by normalization and activation layers. Incorporating these intermediate layers allowed the convolutional layers to selectively utilize non-contextual or lower-level features from earlier layers if and when they prove to be more beneficial for the task. This enhanced model achieved a lower MLM validation loss of 1.43, indicating a clear improvement in capturing relevant token-level information. In the final part of the architecture, the [CLS] token from the convolutional chain output was compressed into a 128-length representation and subsequently used in downstream tasks.
As a feasibility study, extraction of image feature vectors were initially made with a simple AutoEncoder, utilizing an EfficientNetV2-Large model as the Encoder and series of convolutional layers as a simple decoder. Utilizing a previous experience with such models, recreation losses were calculated using MS-SSIM since it provided a robustness towards various issues that might occur with MSE (Blurring of target, image scale issues). Since the goal of the development is creating a production ready, searchable and indexable representation vector, outputs of the EfficientNet were compressed to 128, which was decided during trials with randomly generated vectors of different dimensions in order to adhere to the production requirements. This vector size made it possible to scalably index and search available product data.
After conclusion of the feasibility study, research on image representation vectors was made. In order to retain as much of the contextual information as possible Masked AutoEncoders (MAE) surfaced as a suitable approach. Contextual information is found necessary because simple reconstruction mostly focuses on colors and shapes while what the object is also important in the E-commerce domain.
Fundamentally, a MAE’s primary objective is to extract information from an image by examining its individual components. This forces the model to “understand” the objects in the image in order to create a complete reconstruction. As a brief summary to inner workings of MAE therefore can be given as follows;
  • Divide image into N patches,
  • Randomly select %X of N, mask selected,
  • Using an image transformer (DeiT), create features from each unmasked patch,
  • Reconstruct masked patches with the information gained from unmasked patches.
By default, most of the successful MAE algorithms [12], create relatively large representation vectors. While this approach proved successful in recreating images, our requirement for compressed representations made this logic impractical. As a solution, an architecture to create smaller representations was constructed. Specifically, the architecture was designed to work with the DeiT backbone, compressing and concatenating the created per patch representations to a singular 128 length vector.
The resulting representation was used as the output from the “forward” path of the model with a single 128 sized representation vector for each image in the batch. This representation is then fed through convolutional layers to be converted into expected dimensions for MAE decoder. Following this, output was further processed within the decoder, together with the masked token ids, unmasked token ids and unmasked patches (same as in normal MAE) to reconstruct the input image.
While providing a robust starting point for contextual image understanding, MAE has an inherent issue in image feature extraction. Since it randomly selects patches, some small differences within images has a possibility to not be represented in the resulting vector. These sort of differences in objects is a common occurrence in the e-commerce domain, hence retention of these differences becomes relatively important when deciding whether two objects match each other or not. Our internal analysis showed that most of the object details exist on the center of the image. In order to leverage this a more strict patch selection logic was set up. Instead of random patch selection, dataloader was modified to divide the image into 9 equal patches and then create 4 patches as follows;
  • Select center patch, for details that should be preserved,
  • 2 random patches to retain the contextual understanding of MAE,
  • Entire resized image, for details that could help information retention.
Figure 4: Scale differences of a product. (a) image does not have a white border, (b) image has a white border
From this input a full image was then reconstructed. Another variant that selects 3 random patches in addition to the center patch was also trained. However, this variant did not achieve the expected improvement. This resulted in the hypothesis that the overall outline of the object acquired from the resized full image, assists context understanding and image reconstruction.
As opposed to the initial implementation of MAE, since the input patches include the entire image hence did not have the strict “patch ids” required, recreating just the “masked” parts of the image were not immediately feasible. Therefore this iteration of the architecture learned to reconstruct the full image from the compressed vector.
In the matching trials, an issue kept surfacing regarding the object scales in images. Problems occurred when objects did not take the same space within the image, creating empty spaces between objects and the image borders. This caused vectors to be distanced from each other as if they were variants instead of the same product. An example of this phenomenon can be seen in Figure  4.
Initially as a solution, a separate module was added prior to the dataloader as a data preprocessing method to remove “boring” pixels, closing the gap between object and image border. However this approach quickly proved costly for the production environment, given the sheer number of images processed. Instead, in order to increase the robustness of the model, the same boring pixel removal module has been used in the dataloader, tied to a random number generator. Algorithm for which is as follows:
  • If random number is less than the threshold proceed, if not return original image,
  • Convert image to black and white,
  • Find first non-zero pixel coordinates as (x1,y1), (x2,y2), (x3,y3) and (x4,y4),
  • Extract the center box with found coordinates as points on the edge of a bounding box,
  • Create a background image with the colour as the ”boring pixels”,
  • Scale the bounding box randomly,
  • Embed bounding box to the generated background and return as input image.
Figure 5: Input and Recreation Results of the final iteration of Image Model
During training, input image is cleaned from boring pixels and marked as the reconstruction target. If the generated random number is below the threshold, input of the model will be the image generated with the aforementioned algorithm. If not, input image will be the unprocessed original image. Image reconstruction outputs are provided in Figure 5.
In the training procedure, AdamW optimizer was utilized with a starting learning rate of 0.0001 then, a cosine scheduler scaled this value down to near ”0”. Image and text models are trained until convergence. While trials with different optimizers and learning rate values are made, this setup proved most efficient, resulting in lowest loss values.
In the culmination of the study, we address the challenge of product matching in the e-commerce domain by leveraging previously explained multimodal vector data. As detailed in the preceding sections, we employed sophisticated vector extraction techniques to generate dense embeddings from both textual and visual inputs. These vectors serve as robust representations capturing the semantic and visual nuances of each product. Recognizing that product similarity can be ambiguous when relying solely on either text or image data, we hypothesized that combining both modalities would provide a more comprehensive basis for accurate matching. While textual embeddings encapsulate descriptive and categorical information, visual embeddings capture design, color, and style attributes. By integrating these two sources of information, our approach aims to mitigate the limitations inherent in single-modal matching and enhance the overall precision of product pair assessments.
To efficiently handle the high-dimensional vector data and facilitate rapid similarity searches, we utilized Milvus, an open-source vector database optimized for scalable similarity search. Milvus enabled us to perform vector-based retrieval operations with low latency, ensuring that our system remains responsive even as the product catalog expands. By indexing the combined text and image vectors in Milvus, we streamlined the retrieval process, allowing us to quickly identify candidate product pairs for further evaluation by the decider model. In order to evaluate the memory usage of Milvus, we indexed various sizes of randomized vectors into a single-server Milvus setup and measured the system’s memory consumption. This approach allowed us to assess how different vector sizes impacted memory usage. The results, presented in Table  III, indicated that indexing tens of millions of vectors would be practical only with a reduced vector size of 128.
TABLE III: RAM Usage of 10 Million Vectors on Milvus
Vector Size RAM Usage (GB)
128 5.5
256 10.2
512 20
1024 39.7
In addition to vector size analysis, in Table IV the inference speed comparison with well established multimodal feature extractors are shown, all architectures here have been used in evaluation mode with the same GPU and batch size of 10, using a single worker dataloader. As can be seen, the internal model is considerably faster when inferred, which is a crucial component in a deployed environment. In addition to the infer speed, default vector output sizes for both CLIP and BLIP is far larger, making the usage of these models without modification impossible as per previous vector size feasibility results.
Building upon the vector representations, we developed a decider model designed to discern whether two products constitute a match. The model architecture, implemented using PyTorch Lightning, integrates both textual and visual features to make informed classification decisions. Specifically, the “decider” model concatenates the text and image vectors of each product pair, resulting in comprehensive feature vectors that encapsulate information from both modalities.
The architecture comprises a series of convolutional and fully connected layers. A 1D convolutional layer processes the stacked feature pairs, followed by normalization and activation to stabilize training. Subsequent linear layers further reduce the dimensionality, incorporating dropout layers to mitigate overfitting. The final output layer employs a Softmax activation function to produce probability scores for the binary classification task—determining whether a pair of products matches or not.
TABLE IV: Inference Times per 1000 Product Pairs
Model Generation Time (sec)
Ours 10
CLIP 62
BLIP 302
For training the decider model, similar to image and text model we also employed the AdamW optimizer, since it is known to be performant, complemented by a learning rate scheduler that adjusts the learning rate based on validation loss improvements. The model was optimized using cross-entropy loss, aimed at minimizing classification errors. We trained the model over multiple epochs, ensuring convergence and stability in performance. The training and test datasets, consisting of 350K and 8K product pairs respectively, were obtained through an internal labeling process conducted by catalog experts. Due to the absence of a publicly available, independent test set for Turkish e-commerce product matching, our system’s performance was evaluated against an existing, black-box third-party111a commercially available API-based deduplication service solution. While this methodology introduces a known bias—as the labeled data was created from pairs generated by both our and the third-party algorithms—it served as a necessary starting point to identify potential areas for improvement and to demonstrate that our results surpassed the current solution from a business standpoint. Using the acquired test set, we compared the performance of the trained system with a third-party provider by inspecting the results of each system individually. Our system achieved 0.90 macro average F1 score and the third-party system achieved 0.83. General results of both systems can be seen in Table  V. With a better Recall performance on the match class, our system is better at detecting real duplicate products without losing precision over the third-party system.
Throughout our empirical analysis, we observed specific areas where our model exhibited tendencies that warrant further investigation. Notably, the model showed a propensity to classify visually similar but distinct product variants as matches. For instance, products differing only in subtle packaging variations or minor design elements, while fundamentally representing different SKUs, were sometimes grouped together. This suggests that while our multimodal embeddings excel at capturing broad semantic and visual similarities, they occasionally struggle with fine-grained distinctions crucial for differentiating product variants that are visually very close but operationally separate. Addressing these nuances will be a key focus for future refinements.
TABLE V: Performance Comparison Results
Model Class Precision Recall F1
Internal Match 0.91 0.85 0.88
3rd Party Match 0.90 0.73 0.81
Internal Not Match 0.88 0.96 0.92
3rd Party Not Match 0.76 0.94 0.84
6 Conclusion
By strategically combining unsupervised learning with domain-specific adaptations, we developed an efficient and scalable product deduplication method. Compressing vector dimensions while retaining critical semantic information enabled accurate deduplication with reduced storage and computational costs. Our results demonstrate that well-designed compact vectors can effectively capture product listing nuances, facilitating reliable large-scale deduplication.
Integrating multimodal vector representations, combining text and image embeddings, further improved duplicate detection accuracy. Our decider model, leveraging these embeddings, achieved a macro-average F1 score of 0.90, highlighting its robustness. Additionally, the use of Milvus ensures efficient similarity searches in expanding datasets, enhancing operational efficiency and customer experience. Deployed architecture is serving our internal users with 116M222as of August 2025 product vectors daily.
This study underscores the potential of domain-specific machine learning adaptations in e-commerce. Future work will focus on developing a unified multimodal framework that integrates text and image features to enhance matching accuracy. Efforts will also be directed toward creating lighter and faster models to improve scalability in real-world applications. Furthermore, to address issues arising from product variants and the insufficiency of manual attribute controls, we plan to integrate product attribute information into the decider model. Additionally, future work will involve extending the application to other languages. These advancements aim to create a more accurate, efficient, and adaptable product matching system capable of addressing the dynamic challenges of large-scale e-commerce.